# Another look at inference after prediction

### v3

@@ 3 Efficiency analysis
As PB inference methods suppose the availability of at least a small labeled data set, a natural baseline estimator for comparison is $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, which simply regresses $Y$ on ${\bf{X}}$ within $\mathscr{L}$. For a PB inference method to be preferred over this baseline, it should not only yield valid inference, but also improve statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the prediction is informative of $Y$, and perform no worse than estimation based on $\mathscr{L}$ when $\widehat{Y}$ is uninformative of $Y$. For the remainder of this paper, we study the latter property and analyze an estimator very similar to $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ given by
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right),$
where $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ is the solution to $\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$. The only difference between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that all of the data, rather than the unlabeled data, is used to estimate $\bm{\gamma}^{*}$ in the augmentation term (i.e., $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ is replaced by $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$). When $\pi\approx 0$, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}\approx\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ so that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}\approx\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$. In general, it can be shown that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ gains over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when $\widehat{Y}$ is informative of $Y$ (see Appendix B, specifically Section B.3, for a detailed analysis). Additionally, estimators of this form are commonly used in the semi-parametric inference literature, which we draw upon in the subsequent section in our efficiency analysis.
@@ 3.1 Asymptotic analysis of the PPI estimator
To study the asymptotic variance of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, Theorem 1 provides the influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$. All of the asymptotic results for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ are detailed in Appendix A.
@@
where ${\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\ \bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\right]\frac{R_{i}}{\pi}+\left[\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})\right]\frac{\pi-R_{i}}{\pi}\right\}$ and ${\bf{A}}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}$. Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
$\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
We use Theorem 1 to compare the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Corollary 1.1 shows that the benefit of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ scales directly with the predictiveness of $\widehat{Y}$ for $Y$ and inversely with the proportion of labeled data, $\pi$.
@@
The sufficient condition in Corollary 1.1 coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on ${\bf{X}}$. This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al. [2023a] and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix A.4). Here we consider a simple example to demonstrate when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ under homoskedasticity.
@@
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right)\left[2\sigma_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}^{2}\right]{\bf{A}},$
and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if
@@
where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the correlation of the residuals for $Y$ and $\widehat{Y}$ given ${\bf{X}}$. Put another way, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ improves over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with respect to efficiency when $\widehat{Y}$ is a strong predictor for $Y$ after adjusting for ${\bf{X}}$, though $\widehat{Y}$ does not necessarily need to be derived from a ML model that captures the true regression function. The improvement is stronger with increasing correlation of the residuals between $Y$ and $\widehat{Y}$ and decreasing variability in the residuals of $\widehat{Y}$.
@@ 3.2 Asymptotic analysis of the CC estimator
To further analyze the efficiency of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ we draw on the landmark paper that introduced a class of augmented inverse probability weighted (AIPW) estimators which includes all regular asymptotic linear estimators [Robins et al., 1994]. In the current setting, where the goal is to estimate linear regression coefficients $\bm{\beta}^{*}$ with $Y$ partially observed, but $\widehat{Y}$ fully observed, these estimators have influence functions of the form
@@
By Theorem 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is an AIPW estimator and can be seen as an approximation to the optimal estimator with $\bm{\omega}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$. The PPI estimator is therefore not the most efficient estimator in its class and, from Corollary 1.1, can even be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\bm{\Sigma}_{Y,\widehat{Y}}$. A PB inference strategy should enhance statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever possible and, at the very least, should not be negatively impacted by incorporating unlabeled data into estimation. This raises concerns about the practical utility of the PPI estimator.
To improve upon $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, one approach is to assume a parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ that can be estimated from $\mathscr{L}$ for analysis. However, it is often difficult to specify an appropriate parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$, particularly when the relationship between $Y$ and $\widehat{Y}$ is unknown. Moreover, when the model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is misspecified, an efficiency gain is not guaranteed. To overcome these difficulties, Chen and Chen [2000] proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$. Their approach is motivated by the double-sampling design which, apart from differences in terminology (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is analogous to the partially labeled data setting considered here. The CC estimator approximates $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ with
@@
The estimator with this influence function can be written as
@@
obtained with $\mathscr{L}$. Therefore, a simple way to ensure that an estimator like the PPI estimator has asymptotic variance at least as low as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a weight function in front of the augmentation term. In fact, the framework of Robins et al. [1994] implies that all regular asymptotically linear estimators for $\bm{\beta}^{*}$ augment $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with an augmentation term that incorporates information from $\widehat{Y}$. We present the influence function expansion for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in Theorem 2 to further study its asymptotic variance and compare it with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$. All of the asymptotic results provided for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ are detailed in Appendix C.
@@
Following Chen and Chen [2000],
@@
where ${\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y}_{i},{\bf{X}}_{i})\frac{\pi-R_{i}}{\pi}\right]$ and $\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y},{\bf{X}})$ is defined in (1). Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
$\hskip 2.84526pt\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right]\hskip 2.84526pt.$
The subsequent corollaries apply Theorem 2 to show our two key results: (1) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and (2) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ has the highest asymptotic efficiency within the class of estimators with a weighted linear augmentation term. Together these results illustrate that the CC method is both robust to ML model specification and optimizes statistical efficiency, making it preferable to the PPI method in practice.
@@
Let the mean squared error (MSE) of an estimator $\hat{\bm{\beta}}$ for $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$. Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum [2015], page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
From Corollary 2.1, we note that in contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term, $\bm{\Sigma}_{Y,\widehat{Y}}$, in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of $\widehat{Y}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$. This normalization effectively scales the covariance according to the precision of $\widehat{Y}$ and enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Coming back to the example from the previous section, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ possesses this property while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not.

### v4

@@ 3 Efficiency analysis
As PB inference methods suppose the availability of at least a small labeled data set, a natural baseline for comparison is the estimator based on $\mathscr{L}$ only, $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. For a PB inference method to be preferred over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, it should not only yield valid inference, but also achieve gains in statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the prediction is informative of $Y$ and perform no worse when $\widehat{Y}$ is uninformative. For the remainder of this paper, we study the latter property and analyze a modified version of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ given by
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right).$
The only difference between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that all of the data, rather than the unlabeled data, is used to estimate $\bm{\gamma}^{*}$ in the augmentation term (i.e., $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ is replaced by $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$). When $\pi\approx 0$, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}\approx\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ so that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}\approx\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$. In general, it can be shown that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ gains over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when $\widehat{Y}$ is informative of $Y$ (see Appendix, specifically Section B.3, for a detailed analysis). Additionally, estimators of this form are commonly used in the semi-parametric inference literature, which we draw upon in the subsequent section in our asymptotic analysis.
@@ 3.1 Asymptotic analysis of the PPI estimator
To study the asymptotic propeties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we first derive its influence function expansion in Theorem 1. The detailed derivations are deferred to the Appendix.
@@
where ${\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\ \bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\right]\frac{R_{i}}{\pi}+\left[\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})\right]\frac{\pi-R_{i}}{\pi}\right\}$ and ${\bf{A}}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}$. Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance $\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
We use Theorem 1 to compare the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Corollary 1.1 shows that the benefit of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ scales directly with the predictiveness of $\widehat{Y}$ for $Y$ and inversely with the proportion of labeled data $\pi$.
@@
The sufficient condition in Corollary 1.1 coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on ${\bf{X}}$. This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al 21 and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix, Section A.4). Here we consider a simple example to demonstrate when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ under homoskedasticity.
@@
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right)\left[2\sigma_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}^{2}\right]{\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if
@@
where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the correlation of the residuals for $Y$ and $\widehat{Y}$ given ${\bf{X}}$. Thus $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\widehat{Y}$ is a strong predictor for $Y$ after adjusting for ${\bf{X}}$, though $\widehat{Y}$ does not necessarily need to be derived from a ML model that captures the true regression function. The improvement is stronger with increasing correlation of the residuals between $Y$ and $\widehat{Y}$ and decreasing variability in the residuals of $\widehat{Y}$.
@@ 3.2 Asymptotic analysis of the CC estimator
To further analyze the efficiency of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we draw on the landmark paper that introduced a class of augmented inverse probability weighted (AIPW) estimators which includes all regular asymptotic linear estimators. 24 In the current setting, where the goal is to estimate $\bm{\beta}^{*}$ with $Y$ partially observed and $\widehat{Y}$ fully observed, these estimators have influence functions of the form
@@
By Theorem 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is thus an AIPW estimator and can be seen as an approximation to the optimal estimator with $\bm{\omega}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$. The PPI estimator is therefore not the most efficient estimator in its class and, from Corollary 1.1, can even be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\bm{\Sigma}_{Y,\widehat{Y}}$. A PB inference strategy should enhance statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever possible and, at the very least, should not be negatively impacted by incorporating unlabeled data into estimation. This raises concerns about the practical utility of the PPI estimator.
To improve upon $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, one approach is to assume a parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ that can be estimated from $\mathscr{L}$ for analysis. However, it is often difficult to specify an appropriate model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$, particularly when the relationship between $Y$ and $\widehat{Y}$ is unknown. Moreover, when the model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is misspecified, an efficiency gain is not guaranteed. To overcome these difficulties, Chen and Chen 22 proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$. Their approach is motivated by the double-sampling design which, apart from differences in terminology (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is analogous to the partially labeled data setting considered here. The CC estimator approximates $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ with
@@
The corresponding estimator can be written as
@@
obtained with $\mathscr{L}$. Therefore, a simple way to ensure that an estimator with the same structure as the PPI estimator has asymptotic variance no greater than that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a weight function in front of the augmentation term. We present the influence function expansion for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in Theorem 2 to further study its asymptotic variance and compare it with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$.
@@
Following Chen and Chen 22,
@@
where ${\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y}_{i},{\bf{X}}_{i})\frac{\pi-R_{i}}{\pi}\right]$ and $\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y},{\bf{X}})$ is defined in (1). Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance $\hskip 2.84526pt\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right]\hskip 2.84526pt.$
The subsequent corollaries apply Theorem 2 to show our two key results: (1) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and (2) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ has the highest asymptotic efficiency within the class of estimators with a weighted linear augmentation term. These results suggest that the CC method is preferable to the PPI method for gaining efficiency.
@@
Let the mean squared error (MSE) of an estimator $\hat{\bm{\beta}}$ for $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$. Analogously to the well-known result for the best linear predictor (e.g.,Bickel and Doksum,28 page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
From Corollary 2.1, we note that in contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term, $\bm{\Sigma}_{Y,\widehat{Y}}$, in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of $\widehat{Y}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$. This normalization effectively scales the covariance according to the precision of $\widehat{Y}$ and enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Returning to the previous example, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ possesses this property while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not.